OpenStudy (anonymous):

2x - 12 = 3y 5x + 4y = 7 Help me with Gaussian substitution please! Work me through it. I want to understand.

OpenStudy (anonymous):

@precal

OpenStudy (anonymous):

@SolomonZelman @Zale101 @zepdrix

zepdrix (zepdrix):

Hey c: So we're doing the substitution method, not the elimination method?

OpenStudy (anonymous):

My teacher said it's Gaussian subst. where you have to multiply each side by a number then cancel out a pair of numbers.

zepdrix (zepdrix):

Oh ok, we call that Gaussian Elimination, then we follow it up with back substitution. If your teacher is calling it that though, no big deal.

OpenStudy (anonymous):

Nope my mistake, I wrote it wrong in my notes haha

zepdrix (zepdrix):

Oh hehe. \[\Large\rm 2x - 12 = 3y\]\[\Large\rm 5x + 4y = 7\]We need all of our x's and y's on the same side. In equation #1, let's subtract 3y from each side, and add 12 to each side.

zepdrix (zepdrix):

\[\Large\rm 2x - 3y = 12\]\[\Large\rm 5x + 4y = 7\]Understand that first step ok? :o

OpenStudy (anonymous):

Yes! That's the part I couldn't understand was moving the variables around.

OpenStudy (anonymous):

I think I can work it myself. I'll try.

zepdrix (zepdrix):

Hmm ok cool :) Yah give it a shot.

OpenStudy (anonymous):

Is the solution (3, -2) ? @zepdrix

zepdrix (zepdrix):

Ooo yay! Nice job!

OpenStudy (anonymous):

Thank you! I appreciate it very much!